2x+3y=9
3x+4y=5
Solve using substitution

Answer by mananth(16949)

2 x + 3 y = 9 .............1
3 x + 4 y = 5 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 3
-8 x -12 y = -36
9 x 12 y = 15
Add the two equations
1 x = -21
/ 1
x = -21
plug value of x in (1)
2 x + 3 y = 9
-42 + 3 y = 9
3 y = 9 + 42
3 y = 51
y = 17
m.ananth@hotmail.ca